If the sides of a square are lengthened by 8 cm, the area becomes 289 cm^2. Find the length of a side of the original square.
The length of a side of the original square is?

Answer by 20071436(1)

the original length of the square side's is 9cm.
u may reach this solution by considering the original side as x (variable)
area of the square is s^2 where s is the side length = x+8
then (x+8)^2=289
taking the square root of both sides then x+8=17
then x=17-8=9cm